# Local converse theorems and Langlands parameters

### v1

@@ Abstract.
Let $F$ be a non Archimedean local field, and $G$ be the $F$-points of a connected quasi-split reductive group defined over $F$. In this paper we propose a converse theorem statement for generic Langlands parameters of $G$ when the Langlands dual group of $G$ is acceptable. We then prove it when $G$ is $F$-split. Then we consider a variant which we prove for $G=\mathrm{G}_{2}(F)$ and all quasi-split classical groups. When $F$ has characteristic zero and assuming the validity of the Gross-Prasad and Rallis conjecture, this latter variant translates via the generic local Langlands correspondence of Jantzen and Liu, into the usual local converse theorems for classical groups expressed in terms of Shahidi’s gamma factors.
@@ 1. Introduction
We now briefly explain the main observations that are made in this paper. We denote by $G$ be the $F$-points of a connected quasi-split reductive group defined over $F$, with Weil-Deligne group ${\mathrm{WD}}_{F}$ and L-group ${}^{L}G$. First, thanks to the Local Langlands correspondence for ${\mathrm{GL}}_{n}(F)$ ([HT] and [Hen3]), the local converse theorem of Henniart or its sharp refinement implies that if $\phi:{\mathrm{WD}}_{F}\to{}^{L}G$ is any admissible homomorphism with associated Weil-Deligne parameter $(\varphi,N)$ (see Section 2.2), and $r:{}^{L}G\to{\mathrm{GL}}(V)$ is an irreducible representation of ${}^{L}G$, then the twisted gamma factors of $r\circ\phi$ determine the ${\mathrm{GL}}(V)$-conjugacy class of $r\circ\varphi$ (see Section 3). Suppose that this holds for any irreducible $r$, then a classical result of Steinberg implies, at least when $G$ is $F$-split, that the conjugacy class of $\varphi(\gamma)$ by the Langlands dual group $\widehat{G}$ is uniquely determined for any $\gamma\in\mathrm{W}_{F}$. Under the assumption that $\widehat{G}$ is acceptable, using in particular results of Yu on acceptability (Section 4) and of Silberger and Zink on the Langlands classification of parameters (Section 5), we prove in Theorem 5.3 that this property in turn implies that the equivalence class of $\varphi$ is uniquely determined. If $\phi$ is moreover assumed to be generic this determines the class of $\phi$ uniquely, this folklore fact having recently been proved rigorously in [DHKM] as we recall in Section 2.2. This leads to our main result Theorem 6.3, to which we refer for more complete statements. Using similar ideas, we prove the variant which coincides with the known local converse theorems via the generic local Langlands correspondence in Theorem 7.2.
@@ Acknowledgement.
We thank an anonymous referee for encouraging us to meditate on the notion of local converse theorem. We warmly thank Guy Henniart, Rob Kurinczuck and Wee-Teck Gan for very useful correspondence. We especially thank Ahmed Moussaoui for answering many questions about this work, and Eyal Kaplan for very useful correspondence and detailed comments on an earlier draft, allowing significant improvement of its exposition. Finally and most importantly, we thank Dipendra Prasad for many useful answers and bringing the notion of acceptability to our attention, as well as for spotting a mistake in a previous version, the correction of which allowed to strengthen our results.
@@ 2. Preliminaries
All representations that we consider have complex coefficients. The main reference for the material in this section are [Tat] and [Bor], and [DHKM, Section 6]. Let $F$ be a non Archimedean local field with residual cardinality $q$, Weil group $\mathrm{W}_{F}$ and Weil-Deligne group ${\mathrm{WD}}_{F}=\mathrm{W}_{F}\times{\mathrm{SL}}_{2}({\mathbb{C}})$. We do not impose any restriction on the characteristic of $F$. We denote by ${\mathrm{Fr}}_{F}$ a geometric Frobenius element of $\mathrm{W}_{F}$ and by $\nu_{F}:\mathrm{W}_{F}\to{\mathbb{C}}^{\times}$ the unique unramified character such that $\nu_{F}({\mathrm{Fr}}_{F})=q^{-1}$. For $\gamma\in\mathrm{W}_{F}$ we denote by $d(\gamma)$ the integer such that $\nu(\gamma)=q^{-d(\gamma)}$. Let $G$ be the group of $F$-points of a quasi-split connected reductive group defined over a local field $F$. We denote by $\widehat{G}$ its Langlands dual group and its $L$-group by ${}^{L}G:=\widehat{G}\rtimes\mathrm{W}_{F}$, where $\mathrm{W}_{F}$ acts trivially on $\widehat{G}$ if $G$ is split over $F$. For example ${}^{L}{\mathrm{GL}}_{n}(F)={\mathrm{GL}}_{n}({\mathbb{C}})\times\mathrm{W}_{F}$. By definition, what we call a representation of ${}^{L}G$ is a representation $r$ which is algebraic on $\widehat{G}$, and such that $r_{|\mathrm{W}_{F}}$ factors through a finite dimensional representation of the Galois group of a finite extension of $F$ splitting $G$.
@@ 2.1. Langlands parameters and their local constants
We say that a continuous homomorphism $\varphi:\mathrm{W}_{F}\to{}^{L}G$ is semi-simple if the projection of its image on $\widehat{G}$ consists of semi-simple elements. We say that a homomorphism $\phi:{\mathrm{WD}}_{F}\to{}^{L}G$ is admissible if $\phi_{|W_{F}}$ is semi-simple, and $\phi_{|{\mathrm{SL}}_{2}({\mathbb{C}})}$ is algebraic. We say that two admissible homomorphisms $\phi_{i}:{\mathrm{WD}}_{F}\to{}^{L}G$ for $i=1,2$ are equivalent, which we write $\phi_{1}\sim\phi_{2}$, if they are conjugate under $\widehat{G}$. We put the same equivalence relations on semi-simple homomorphisms from $\mathrm{W}_{F}$ to ${}^{L}G$. An admissible homomorphism $\phi:{\mathrm{WD}}_{F}\to{}^{L}{\mathrm{GL}}_{n}(F)$ identifies with a (smooth) semi-simple representation of ${\mathrm{WD}}_{F}$ of dimension $n$, because composition with the projection onto ${\mathrm{GL}}_{n}({\mathbb{C}})$ is injective, and we will do this identification freely.
@@
$\mathrm{Ad}(\tau(\gamma))(N)=\nu(\gamma)N$
for all $\gamma\in\mathrm{W}_{F}$. There is a natural action of $\widehat{G}$ on the set of Weil-Deligne parameters given by the formula
@@
and we say that two Weil-Deligne parameters are equivlent if they are in the same $\widehat{G}$-orbit. Let $\phi:{\mathrm{WD}}_{F}\to{}^{L}G$ be an admissible homomorphism, and denote by $d\phi$ the differential of $\phi_{|{\mathrm{SL}}_{2}({\mathbb{C}})}$. As explained in [DHKM, Remark 6.4, (2)], one asociates to $\phi$ the Weil-Deligne parameter $(\varphi,N)$ where
@@
$\gamma(s,\phi,\psi)=\gamma(s,\varphi,\psi).$
@@ 2.2. Generic parameters
Let $\phi_{1},\phi_{2}:{\mathrm{WD}}_{F}\to{}^{L}G$ be two generic homomorphisms, then $\phi_{1}\sim\phi_{2}\iff\varphi_{1}\sim\varphi_{2}$.
@@
This follows from the equivalence between (1) and (2) in [DHKM, Proposition 6.10]. ∎
@@ 3. A consequence of the sharp local converse theorem for ${\mathrm{GL}}_{n}$
$|\det|_{F}^{(1-n)/2}\rho\times\dots\times|\det|_{F}^{(n-1)/2}\rho.$
@@
$\mathrm{St}_{r_{1}}(\rho_{1})\times\dots\times\mathrm{St}_{r_{1}}(\rho_{r})$
@@
${\mathrm{LQ}}(\mathrm{St}_{r_{1}}(\rho_{1})\times\dots\times\mathrm{St}_{r_{1}}(\rho_{r})).$
@@
$\mathrm{LLC}(\sum_{i=1}^{t}\varphi_{i}\otimes{\mathrm{Sp}}(r_{i}))={\mathrm{LQ}}(\mathrm{St}_{r_{1}}(\rho_{1})\times\dots\times\mathrm{St}_{r_{1}}(\rho_{r})).$
@@
Write $\pi$ as ${\mathrm{LQ}}(\mathrm{St}_{r_{1}}(\rho_{1})\times\dots\times\mathrm{St}_{r_{1}}(\rho_{r}))$ as in Lemma 3.1. Then its Langlands parameter is $\phi:=\sum_{i=1}^{t}\varphi_{i}\otimes{\mathrm{Sp}}(r_{i})$ where $\mathrm{LLC}(\varphi_{i})=\rho_{i}$. First, this implies that the restriction to $\mathrm{W}_{F}$ of $\phi$ is $\oplus_{i=1}^{t}(\varphi_{i}\oplus\dots\oplus\varphi_{i})$. Now set
@@
The direct implication follows from the multiplicativity properties of gamma factors. Let us explain the converse direction. Let $\pi_{1}$ be the irreducible representation of ${\mathrm{GL}}_{n}(F)$ parametrized by $\phi_{1}$, and $\pi_{2}$ that parametrized by $\phi_{2}$. We denote by $\tau_{1}$ and $\tau_{2}$ respectively, the unique generic representation of ${\mathrm{GL}}_{n}(F)$ with same cuspidal support as $\pi_{1}$ and $\pi_{2}$. Thanks to the preservation of local constant properties of the local Langlands correspondence, and by multiplicativity of Rankin-Selberg gamma factors ([JPSS1]), we have
@@ 4. Acceptable reductive algebraic groups over ${\mathbb{C}}$
First for a general group $H$ and $X$ a subset of $H$, we denote by $C_{H}(X)$ the centralizerof $X$ in $H$. Let $\widehat{G}$ be the group of complex points of a connected reductive algebraic group. Following [Lar1] and [Yu] we define the notion of (strong) acceptability.
@@
The notion of acceptability is now well understood, and we refer for example to [Lar1], [Lar2], [CG] and [Yu]. In particular, according to [Lar1, Proposition 1.7] and [Yu, Theorem 1.1], we have the following result.
@@
For the first part, we let $\theta$ be a Cartan involution of $\widehat{G}$. Up to conjugating $\widehat{L}$, we may assume that it is stabilized by $\theta$. Denote by $U$ the connected center of $\widehat{L}^{\theta}$, in this situation we have $\widehat{L}^{\theta}=C_{\widehat{G}^{\theta}}(U)$ and $\widehat{L}^{\theta}$ is known to be connected (see the references in [Yu, Section 2]). By Theorem 4.2 the group $\widehat{G}^{\theta}$ is strongly acceptable, hence so is $\widehat{L}^{\theta}$ thanks to [Yu, Lemma 2.3]. Finally we deduce that $\widehat{L}$ is acceptable thanks to Theorem 4.2 again. For the second part, both $\varphi$ and $\varphi^{\prime}$ have their respective image in some possibly different maximal compact subgroup of $\widehat{G}$. Because such maximal compact subgroups are all conjugate to one another, we may assume that $\varphi$ and $\varphi^{\prime}$ actually have their image in $\widehat{G}^{\theta}$. The result now follows from the strong acceptability of $\widehat{G}^{\theta}$, stated in Theorem 4.2. ∎
@@ 5. A consequence of results of Silberger and Zink
Then by [SZ, Propositions 5.5] we have the following result.
@@
By local conjugacy, we can suppose that $z(\varphi)=z(\varphi^{\prime})$. Now take $\gamma\in W_{F}$, we have $\varphi^{\prime}(\gamma)=x_{\gamma}\varphi(\gamma)x_{\gamma}^{-1}$ for some $x_{\gamma}\in\widehat{G}$. By uniqueness of the elliptic and hyperbolic part of a semi-simple element, this implies that $x_{\gamma}$ centralizes $z(\varphi)$, i.e. $x_{\gamma}\in\widehat{L}_{\varphi}$. It also implies that $\varphi^{\prime}_{t}(\gamma)$ and $\varphi_{t}(\gamma)$ are conjugate by $x_{\gamma}$. In other words $\varphi^{\prime}_{t},\varphi^{\prime}_{t}:\mathrm{W}_{F}\to\widehat{L}_{\varphi}$ are two homomorphisms with bounded image, locally $\widehat{L}_{\varphi}$-conjugate. By Proposition 4.3 they are globally $\widehat{L}_{\varphi}$-conjugate. This in turns implies that $\varphi$ and $\varphi^{\prime}$ are globally conjugate (actually under $\widehat{L}_{\varphi}$). ∎

### v4

@@ Abstract.
Let $F$ be a non Archimedean local field, and $G$ be the $F$-points of a connected quasi-split reductive group defined over $F$. In this paper we prove a converse theorem statement for generic Langlands parameters of $G$ when $G$ is $F$-split and the Langlands dual group of $G$ is acceptable, and propose a possible extension of it without the split assumption. We also prove that the statement does not apply to ${\mathrm{SO}}_{6}(F)$ for certain choices of $F$. Then we consider a variant which we prove for $G=\mathrm{G}_{2}(F)$ and all quasi-split classical groups. When $F$ has characteristic zero and assuming the validity of the Gross-Prasad and Rallis conjecture, this latter variant translates via the generic local Langlands correspondence of Jantzen and Liu, into the usual local converse theorems for classical groups expressed in terms of Shahidi’s gamma factors.
@@ 1. Introduction
We now briefly explain the main observations that are made in this paper. We denote by $G$ be the $F$-points of a connected quasi-split reductive group defined over $F$, with Weil-Deligne group ${\mathrm{WD}}_{F}$ and L-group ${}^{L}G$. First, thanks to the Local Langlands correspondence for ${\mathrm{GL}}_{n}(F)$ ([HT] and [Hen3]), the local converse theorem of Henniart or its sharp refinement implies that if $\phi:{\mathrm{WD}}_{F}\to{}^{L}G$ is any admissible homomorphism with associated Weil-Deligne parameter $(\varphi,N)$ (see Section 2.2), and $r:{}^{L}G\to{\mathrm{GL}}(V)$ is an irreducible representation of ${}^{L}G$, then the twisted gamma factors of $r\circ\phi$ determine the ${\mathrm{GL}}(V)$-conjugacy class of $r\circ\varphi$ (see Section 3). Suppose that this holds for any irreducible $r$, then, when $G$ is $F$-split, a classical result of Steinberg implies that the conjugacy class of $\varphi(\gamma)$ by the Langlands dual group $\widehat{G}$ is uniquely determined for any $\gamma\in\mathrm{W}_{F}$. Under the assumption that $\widehat{G}$ is acceptable, using in particular results of Yu on acceptability (Section 4) and of Silberger and Zink on the Langlands classification of parameters (Section 5), we prove in Theorem 5.3 that this property in turn implies that the equivalence class of $\varphi$ is uniquely determined. If $\phi$ is moreover assumed to be generic this determines the class of $\phi$ uniquely, this folklore fact having recently been proved rigorously in [DHKM] as we recall in Section 2.2. This leads to our main result Theorem 6.3, to which we refer for more complete statements. We also prove that the statement of Theorem 6.3 does not apply to the group ${\mathrm{SO}}_{6}(F)$ for some specific fields $F$, by studying unacceptable parameters in Section 7. Finally using ideas similar to those of our main theorem, we prove the variant which coincides with the known local converse theorems via the generic local Langlands correspondence in Theorem 8.2.
@@ Acknowledgement.
We thank an anonymous referee for encouraging us to meditate on the notion of local converse theorem. We warmly thank Guy Henniart, Wee-Teck Gan and Rob Kurinczuk for very useful correspondence. We especially thank Ahmed Moussaoui for answering many questions about this work, and Eyal Kaplan for very useful correspondence and detailed comments on an earlier draft, allowing significant improvement of its exposition. Finally and most importantly, we thank Dipendra Prasad for many useful answers and bringing the notion of acceptability to our attention, as well as for spotting a mistake in a previous version, the correction of which allowed to strengthen our results.
@@ 2. Preliminaries
All representations that we consider have complex coefficients. The main reference for the material in this section are [Tat] and [Bor], and [DHKM, Section 6]. Let $F$ be a non Archimedean local field with residual cardinality $q$, Weil group $\mathrm{W}_{F}$ and Weil-Deligne group ${\mathrm{WD}}_{F}=\mathrm{W}_{F}\times{\mathrm{SL}}_{2}({\mathbb{C}})$. We do not impose any restriction on the characteristic of $F$. We denote by $I_{F}$ the inertia subgroup of $\mathrm{W}_{F}$, and fix ${\mathrm{Fr}}_{F}$ a geometric Frobenius element of $\mathrm{W}_{F}$. We set $\nu_{F}:\mathrm{W}_{F}\to{\mathbb{C}}^{\times}$ to be the unique character trivial on $I_{F}$ such that $\nu_{F}({\mathrm{Fr}}_{F})=q^{-1}$. Via local class field theory, the map $\nu$ corresponds to the normalized absolute value |||\ || | on $F^{\times}$. For $\gamma\in\mathrm{W}_{F}$ we denote by $d(\gamma)$ the integer such that $\nu(\gamma)=q^{-d(\gamma)}$. Let $G$ be the group of $F$-points of a quasi-split connected reductive group defined over a local field $F$. We denote by $\widehat{G}$ its Langlands dual group and its $L$-group by ${}^{L}G:=\widehat{G}\rtimes\mathrm{W}_{F}$, where $\mathrm{W}_{F}$ acts trivially on $\widehat{G}$ if $G$ is split over $F$. For example ${}^{L}{\mathrm{GL}}_{n}(F)={\mathrm{GL}}_{n}({\mathbb{C}})\times\mathrm{W}_{F}$. By definition, what we call a representation of ${}^{L}G$ is a finite dimensional semi-simple and continuous representation.
@@ 2.1. Langlands parameters and their local constants
[…]
@@
(2.1) $\mathrm{Ad}(\tau(\gamma))(N)=\nu(\gamma)N$
for all $\gamma\in\mathrm{W}_{F}$. We denote by $\mathcal{N}_{\tau}$ the space of (necessary nilpotent) elements in $\widehat{\mathfrak{g}}$ satisfying Equation (2.1), and by $\widehat{G}_{\tau}$ the centralizer of the image of $\tau$ in $\widehat{G}$. There is a natural action of $\widehat{G}$ on the set of Weil-Deligne parameters given by the formula
@@
and we say that two Weil-Deligne parameters are equivlent if they are in the same $\widehat{G}$-orbit. Let $\phi:{\mathrm{WD}}_{F}\to{}^{L}G$ be an admissible homomorphism, and denote by $d\phi$ the differential of $\phi_{|{\mathrm{SL}}_{2}({\mathbb{C}})}$ at $I_{2}$. As explained in [DHKM, Remark 6.4, (2)], one asociates to $\phi$ the Weil-Deligne parameter $(\varphi,N)$ where
@@
(2.3) $\gamma(s,\phi,\psi)=\gamma(s,\varphi,\psi).$
@@ 2.2. Generic parameters
[…]
@@
This follows from the equivalence between (1) and (2) in [DHKM, Proposition 6.10], which asserts that an admissible homomorphism $\phi:{\mathrm{WD}}_{F}\to{}^{L}G$ is generic if and only if its corresponding Weil-Deligne parameter $(\varphi,N)$ is such that $N$ lies in the unique open $\widehat{G}_{\varphi}$-orbit in $\mathcal{N}_{\varphi}$. ∎
@@ 3. A consequence of the sharp local converse theorem for ${\mathrm{GL}}_{n}$
$|\det|_{F}^{(1-r)/2}\rho\times\dots\times|\det|_{F}^{(r-1)/2}\rho.$
@@
$\mathrm{St}_{r_{1}}(\rho_{1})\times\dots\times\mathrm{St}_{r_{t}}(\rho_{t})$
@@
${\mathrm{LQ}}(\mathrm{St}_{r_{1}}(\rho_{1})\times\dots\times\mathrm{St}_{r_{t}}(\rho_{t})).$
@@
$\mathrm{LLC}(\sum_{i=1}^{t}\varphi_{i}\otimes{\mathrm{Sp}}(r_{i}))={\mathrm{LQ}}(\mathrm{St}_{r_{1}}(\rho_{1})\times\dots\times\mathrm{St}_{r_{t}}(\rho_{t})).$
@@
Write $\pi$ as ${\mathrm{LQ}}(\mathrm{St}_{r_{1}}(\rho_{1})\times\dots\times\mathrm{St}_{r_{t}}(\rho_{t}))$ as in Lemma 3.1. Then its Langlands parameter is $\phi:=\sum_{i=1}^{t}\varphi_{i}\otimes{\mathrm{Sp}}(r_{i})$ where $\mathrm{LLC}(\varphi_{i})=\rho_{i}$. First, this implies that the restriction to $\mathrm{W}_{F}$ of $\phi$ is $\oplus_{i=1}^{t}(\varphi_{i}\oplus\dots\oplus\varphi_{i})$. Now set
@@
The direct implication follows from Equation (2.3). Let us explain the converse direction. Let $\pi_{1}$ be the irreducible representation of ${\mathrm{GL}}_{n}(F)$ parametrized by $\phi_{1}$, and $\pi_{2}$ that parametrized by $\phi_{2}$. We denote by $\tau_{1}$ and $\tau_{2}$ respectively, the unique generic representation of ${\mathrm{GL}}_{n}(F)$ with same cuspidal support as $\pi_{1}$ and $\pi_{2}$. Thanks to the preservation of local constant properties of the local Langlands correspondence, and by multiplicativity of Rankin-Selberg gamma factors ([JPSS1]), we have
@@ 4. Acceptable reductive algebraic groups over ${\mathbb{C}}$
First for a general group $H$ and $X$ a subset of $H$, we denote by $C_{H}(X)$ the centralizerof $X$ in $H$. Let $\widehat{G}$ be a connected reductive complex Lie group. Following [Lar1] and [Yu] we define the notion of (strong) acceptability.
@@
The notion of acceptability is now well understood, and we refer for example to [Lar1], [Lar2], [CG] and [Yu]. In particular, according to [Lar1, Propositions 1.6 and 1.7] and [Yu, Theorem 1.1], we have the following result.
@@
    Let $\widehat{K}$ be a maximal compact subgroup of $\widehat{G}$, then two elements of $\widehat{K}$ are $\widehat{K}$-conjugate if and only if they are $\widehat{G}$-conjugate.
  2. (b)
@@
For the first part, we let $\theta$ be a Cartan involution of $\widehat{G}$. Up to conjugating $\widehat{L}$, we may assume that it is stabilized by $\theta$. Denote by $U$ the connected center of $\widehat{L}^{\theta}$, in this situation we have $\widehat{L}^{\theta}=C_{\widehat{G}^{\theta}}(U)$ and $\widehat{L}^{\theta}$ is known to be connected (see the references in [Yu, Section 2]). By Theorem 4.2 the group $\widehat{G}^{\theta}$ is strongly acceptable, hence so is $\widehat{L}^{\theta}$ thanks to [Yu, Lemma 2.3]. Finally we deduce that $\widehat{L}$ is acceptable thanks to Theorem 4.2 again. For the second part, both $\varphi$ and $\varphi^{\prime}$ have their respective image in some possibly different maximal compact subgroup of $\widehat{G}$. Because such maximal compact subgroups are all conjugate to one another, we may assume that $\varphi$ and $\varphi^{\prime}$ actually have their image in $\widehat{G}^{\theta}$. The result now follows from (a) and (b) of Theorem 4.2. ∎
@@ 5. A consequence of results of Silberger and Zink
Then by [SZ, Proposition 5.5] we have the following result.
@@
By local conjugacy, we can suppose that $\varphi({\mathrm{Fr}}_{F})=\varphi^{\prime}({\mathrm{Fr}}_{F})=:f$. This means that $\varphi_{t}({\mathrm{Fr}}_{F})=\varphi_{t}^{\prime}({\mathrm{Fr}}_{F})$ and $z(\varphi)=z(\varphi^{\prime}):=z$. Let $\widehat{L}:=\widehat{L}_{z}$ be the centralizer of $z$ in $\widehat{G}$. Hence the semi-simple homomorphisms $\varphi_{t}$ and $\varphi_{t}^{\prime}$ are $\widehat{L}$ valued, and we want to prove that they are locally $\widehat{L}$-conjugate. If $\gamma\in\mathrm{W}_{F}-I_{F}$, then the $\widehat{G}$-conjugacy of $\varphi(\gamma)=\varphi^{\prime}(\gamma)$ and the uniqueness of the polar decomposition imply that they are $C_{\widehat{G}}(z^{d(\gamma)})$-conjugate, i.e. that they are $\widehat{L}$-conjugate because $C_{\widehat{G}}(z^{d})=C_{\widehat{G}}(z)$ for any $d\neq 0$ by hyperbolicity of $z$. If $\gamma=\iota\in I_{F}$, the argument is a refinement of the previous one. First one checks that there exists $d\in{\mathbb{N}}^{*}$ such that:
  1. (a)
    $f^{d}$ commutes with the finite groups $\varphi(I_{F})$ and $\varphi^{\prime}(I_{F})$,
  2. (b)
    for any $k\geq 1$, $C_{\widehat{G}}(f^{dk})=C_{\widehat{G}}(f^{d})$.
Now there exists $g\in\widehat{G}$ such that $g\varphi({\mathrm{Fr}}_{F}^{d}\iota)g^{-1}=\varphi^{\prime}({\mathrm{Fr}}_{F}^{d}\iota)$ by assumption. By considering hyperbolic parts, we deduce that $g=l\in C_{\widehat{G}}(z^{d})=\widehat{L}$. Moreover exponentiating te above equality to some power $k\geq 1$ such that $\varphi(\iota)^{k}=\varphi^{\prime}(\iota)^{k}=1$, we deduce that $g\in C_{\widehat{G}}(f^{dk})=C_{\widehat{G}}(f^{d})$. From this we deduce that $l\varphi(\iota)l^{-1}=\varphi^{\prime}(\iota)$. Hence $\varphi_{t}$ and $\varphi_{t}^{\prime}$ are locally $\widehat{L}$-conjugate and bounded, so they are globally $\widehat{L}$-conjugate thanks to Proposition 4.3. This implies that $\varphi$ and $\varphi^{\prime}$ are globally conjugate. ∎